how can I know how many gallions of paints to cover my living room that measures 30ft by 18ft?
Answer by ankor@dixie-net.com(22740)

how can I know how many gallons of paints to cover my living room that measures 30ft by 18ft?
:
Find the area of the painted area. If you are covering all 4 walls and the ceiling is 8 ft high:
2(30*8) + 2(18*8) =
2(240) + 2(144) = 768 sq/ft,
if you include the ceiling, add another 30*18 = 540 sq/ft
:
Of course it will probably be a little less because of doors and windows, but this will give you a good idea
:
Paint cans tell you how many sq ft it will cover, usually around 450 sq/ft
Divide the total area you are going to paint by that amt, round it up to the nearest gallon.
:
If you are changing the color drastically you may need to paint it twice.